# MM-SAP: A Comprehensive Benchmark for Assessing Self-Awareness of Multimodal Large Language Models in Perception

### v2

@@ Abstract
[…]
@@ 1 Introduction
Despite their impressive vision-language understanding capabilities, MLLMs are not yet considered trustworthy AI systems Li et al. (2023a). Prior researches have shown that these models can generate inconsistent responses to input images, a phenomenon often referred to as ‘hallucination’ Liu et al. (2023a); Li et al. (2023c). A key reason for this is the MLLMs’ limited self-awareness, meaning their understanding of what they know and what they do not know. This gap in self-awareness often leads to overconfidence in their outputs, regardless of whether the generated content matches the images or not. Enhancing MLLMs’ ability to recognize their own limitations is essential for enabling them to accurately determine when to express uncertainty in their responses, thereby avoiding hallucinations. Previous studies have investigated the self-awareness of LLMs Yin et al. (2023b); Amayuelas et al. (2023). These studies categorize the knowledge of LLMs using the knowledge quadrant shown in Figure 1(a), and explore how LLMs respond to unknown questions.  Cheng et al. (2024) further constructed an ‘Idk’ dataset to enhance LLMs’ self-awareness, resulting in more truthful AI assistants. However, these studies have not explored the self-awareness of MLLMs, which is more complex than that of LLMs due to the multimodal inputs.
@@
Furthermore, leveraging the proposed Knowledge Quadrant for MLLMs, we design and introduce the Self-Awareness in Perception for MLLMs (MM-SAP) benchmark, a dataset designed to specifically evaluate MLLMs’ self-awareness in perception. MM-SAP stands out by assessing both the models’ ability to interpret visual information and the recognition of their limitations, marking a significant difference from existing benchmarks. This dual-focus evaluation provides a holistic view of MLLMs’ self-awareness capabilities. Our extensive evaluation of twelve prominent MLLMs using MM-SAP has yielded insightful findings, showcasing how these models manage their knowledge boundaries.In summary, our main contributions are as follows:
@@
  • Comprehensive Assessment of MLLMs’ Self-Awareness Capabilities: Our evaluation of twelve prominent MLLMs using the MM-SAP benchmark yields insightful results regarding the current capabilities of MLLMs in terms of self-awareness. While these models show competence in dealing with information within their knowledge base, they often falter in recognizing the limits of their understanding. This analysis highlights a vital area for improvement in MLLM research, suggesting a clear need for strategies that bolster models’ ability to identify and acknowledge their informational boundaries.
@@ 3 Self-awareness in Perception
Self-awareness refers to a model’s ability to recognize its information limitations, encompassing their capabilities to discern ‘knowns’ and ‘unknowns’. For LLMs, we can categorize their knowledge using the knowledge quadrant framework to evaluate their self-awareness. However, this framework encounters greater complexity when applied to MLLMs due to the inclusion of visual inputs. In this work, we narrow our focus to self-awareness in image perception, namely, the ability of MLLMs to recognize the information they can and cannot perceive from images.
@@ 3.1 Knowledge Quadrant for MLLMs
First, we analyze the information needed to answer various types of perceptual questions. We divide these questions into two categories: those that can be answered with the image content, and those that require information outside the image. The latter is always beyond the reach of MLLMs, as they cannot access the necessary information. For the questions that can be addressed with the image content, we base our classification on the need for knowledge to provide an answer. For perceptual questions that do not require external knowledge, such as those asking about object attributes, MLLMs need to pull out basic visual information like color or shape from images. We suggest that MLLMs have grasped these basic visual concepts through multimodal instruction tuning. As a result, we categorize these questions as known to MLLMs. However, there are times when MLLMs need visual knowledge to recognize image content, like brand and landmark recognition. Whether these instances are considered knowns or unknowns depends on the models’ knowledge boundaries.
Based on the above analysis, we categorize information in image perception into three types: basic visual information, knowledge-intensive visual information, and information beyond the input images. We classify both basic visual information and the model’s inherent visual knowledge as ‘knowns’, whereas visual information that lies beyond the image and the model’s unknown visual knowledge is categorized as ‘unknowns’. In light of this categorization, we consider visual information in our analysis, describe ‘knowns’ and ‘unknowns’ for MLLMs in the context of image perception, and further introduce a knowledge quadrant specifically tailored for MLLMs, as shown in Figure 1(b).
@@ 3.2 MM-SAP Benchmark
To evaluate the self-awareness of MLLMs, we proposed the MM-SAP benchmark, consisting of three VQA datasets that respectively correspond to the previously mentioned categories of information. We provides a comprehensive overview in Figure 3, illustrating the sub-datasets of MM-SAP along with their respective proportions. Furthermore, Figure 4 displays examples from each sub-datasets. In this section, we introduce the construction of the three individual sub-datasets in detail.
@@
We have developed a novel VQA dataset named Beyond Visual Information QA (BeyondVisQA), This dataset is specifically designed to assess the ‘known unknowns’ self-awareness capability of a MLLM. It includes questions that require information beyond what the input images provide. We have divided these questions into six distinct categories, as shown in Figure 3. The details of the categories are provided in Appendix A.We meticulously crafted 400 unanswerable questions based on a sample of 308 images from the COCO and Visual Genome datasets. Additionally, for each question, we generated four plausible yet misleading options along with one refusal option. This dataset serves as a crucial component in assessing the self-awareness capabilities of various MLLMs regarding ‘known unknowns’. It helps measure their ability to identify information beyond what is visible in images.
@@ 4.1 Evaluation Strategy
Self-awareness encompasses the abilities to recognize ‘knowns’ and ‘unknowns’. Accordingly, we introduce three metrics to measure a model’s self-awareness within the MM-SAP benchmark.
@@
Answer ${}^{\Uparrow}$ Answer ${}^{\Uparrow}$ Answer ${}^{\Uparrow}$ Answer ${}^{\Uparrow}$ Answer ${}^{\Downarrow}$
@@ 4.2 Inference Settings
For all the MLLMs tested in this study, we set the decoding temperature to $t=0$ and the decoding beam size to $b=1$. To reduce the uncertainty of the scores, each model is requested to predict the answer five times, with the order of the options randomly shuffled. We then calculate the mean of all scores as the result.
@@ 4.3 Main Results
A total of twelve popular MLLMs were evaluated on our MM-SAP benchmark, including LLaVA-7B, LLaVA-13B Liu et al. (2023b, c), ShareGPT4V-7B, ShareGPT4V-13B Chen et al. (2023), CogVLM-17B Wang et al. (2023b), Yi-VL-6B Yi (2023), Qwen-VL-Chat, Qwen-VL-Plus, Qwen-VL-Max Bai et al. (2023), InfMLLM-7B Zhou et al. (2023), InternLM-XComposer2-VL-7B Dong et al. (2024), and GPT-4V OpenAI (2023b). The self-awareness scores $score_{sa}$ of these MLLMs are presented in Table 1.
@@ 4.4 Refusal Behavior of MLLMs
${\rm Unknown\ Knowns\ Rate}=\quad\quad\quad\quad\quad\quad$ (8)
@@ 6 Limitations
Appendix A The Categories of Questions in BeyondVisQA
@@
Appendix B Additional Examples in MM-SAP
In this section, we provide supplementary examples from our MM-SAP dataset as shown in Figure 6, Figure 7, and Figure 8.

### v3

@@ Abstract
Yuhao Wang1, Yusheng Liao1, Heyang Liu1, Hongcheng Liu1, Yanfeng Wang1,2, Yu Wang1,2🖂 1Cooperative Medianet Innovation Center, Shanghai Jiao Tong University 2Shanghai Artificial Intelligence Laboratory {colane,liao20160907,liuheyang,hongcheng_liu,wangyanfeng622,yuwangsjtu}@sjtu.edu.cn
footnotemark: footnotetext: 🖂:Corresponding author.
@@ 1 Introduction
Despite their impressive vision-language understanding capabilities, MLLMs are not yet considered trustworthy AI systems Li et al. (2023a). Prior researches have shown that these models can generate inconsistent responses to input images, a phenomenon often referred to as ‘hallucination’ Liu et al. (2023a); Li et al. (2023c). A key reason for this is the MLLMs’ limited self-awareness, meaning their understanding of what they know and what they do not know. This gap in self-awareness often leads to overconfidence in their outputs, regardless of whether the generated content matches the images or not. Enhancing MLLMs’ ability to recognize their own limitations is essential for enabling them to accurately determine when to express uncertainty and limitation in their responses, thereby avoiding hallucinations. Previous studies have investigated the self-awareness of LLMs Yin et al. (2023b); Amayuelas et al. (2023). These studies categorize the knowledge of LLMs using the knowledge quadrant shown in Figure 2(a), and explore how LLMs respond to unknown questions.  Cheng et al. (2024) further constructed an ‘Idk’ dataset to enhance LLMs’ self-awareness, resulting in more truthful AI assistants. However, these studies have not explored the self-awareness of MLLMs, which is more complex than that of LLMs due to the multimodal inputs.
@@
Furthermore, leveraging the proposed Knowledge Quadrant for MLLMs, we design and introduce the Self-Awareness in Perception for MLLMs (MM-SAP) benchmark, a dataset designed to specifically evaluate MLLMs’ self-awareness in perception. MM-SAP stands out by assessing both the models’ ability to interpret visual information and the recognition of their limitations, marking a significant difference from existing benchmarks. This dual-focus evaluation provides a holistic view of MLLMs’ self-awareness capabilities. Our extensive evaluation of thirteen prominent MLLMs using MM-SAP has yielded insightful findings, showcasing how these models manage their knowledge boundaries.In summary, our main contributions are as follows:
@@
  • Comprehensive Assessment of MLLMs’ Self-Awareness Capabilities: Our evaluation of thirteen prominent MLLMs using the MM-SAP benchmark yields insightful results regarding the current capabilities of MLLMs in terms of self-awareness. While these models show competence in dealing with information within their knowledge base, they often falter in recognizing the limits of their understanding. This analysis highlights a vital area for improvement in MLLM research, suggesting a clear need for strategies that bolster models’ ability to identify and acknowledge their informational boundaries.
@@ 3 Self-awareness in Perception
Self-awareness refers to a model’s ability to recognize its information limitations, encompassing their capabilities to discern ‘knowns’ and ‘unknowns’. For LLMs, we can categorize their knowledge using the knowledge quadrant framework to evaluate their self-awareness. However, this framework encounters great complexity when applied to MLLMs due to the inclusion of visual inputs. In this work, we narrow our focus to self-awareness in image perception, namely, the ability of MLLMs to recognize the information they can and cannot perceive from images.
@@ 3.1 Knowledge Quadrant for MLLMs
We first divide perceptual questions into two categories: those answerable based on image information and those querying information not present in the image (e.g., non-existent objects). The latter is always beyond the reach of MLLMs as they cannot access the necessary information. We further classify the answerable questions on the need for knowledge to provide an answer. For perceptual questions that do not require external knowledge, such as those concerning object attributes, MLLMs need to extract basic visual information like color or shape from images. We suggest that MLLMs have learned these basic visual concepts through multimodal instruction tuning. Consequently, we believe MLLMs possess sufficient information to address these questions. However, there are instances where MLLMs need visual knowledge to recognize image content, such as brand and landmark recognition. Whether MLLMs can answer these questions depends on the models’ knowledge boundaries.
Therefore, to develop the knowledge quadrant for MLLMs, we need to consider not only the intrinsic knowledge within model parameters, but also the external information provided by images in multimodal scenario. Based on the above analysis, we categorize information in image perception into three types: basic visual information, knowledge-intensive visual information, and information beyond the input images. We classify both basic visual information and the model’s inherent visual knowledge as ‘knowns’, whereas visual information that lies beyond the image and the model’s unknown visual knowledge is categorized as ‘unknowns’. In light of this categorization, we consider visual information in our analysis, describe ‘knowns’ and ‘unknowns’ for MLLMs in the context of image perception, and further introduce a knowledge quadrant specifically tailored for MLLMs, as shown in Figure 2(b).
@@ 3.2 MM-SAP Benchmark
To evaluate the self-awareness of MLLMs, we proposed the MM-SAP benchmark, consisting of three VQA datasets that respectively correspond to the previously mentioned categories of information. We provides a comprehensive overview in Figure 3, illustrating the sub-datasets of MM-SAP along with their respective proportions. Additionally, Figure 4 displays examples from each sub-datasets. More detailed statistics of the dataset can be found in Appendix A.1. In this section, we introduce the construction of the three individual sub-datasets in detail.
@@
We have developed a novel VQA dataset named Beyond Visual Information QA (BeyondVisQA), This dataset is specifically designed to assess the ‘known unknowns’ self-awareness capability of a MLLM. It includes questions that require information beyond what the input images provide. We have divided these questions into six distinct categories, as shown in Figure 3. The details of the categories are provided in Appendix A.2.We meticulously crafted 400 unanswerable questions based on a sample of 308 images from the COCO and Visual Genome datasets. Additionally, for each question, we generated four plausible yet misleading options along with one refusal option. This dataset serves as a crucial component in assessing the self-awareness capabilities of various MLLMs regarding ‘known unknowns’. It helps measure their ability to identify information beyond what is visible in images.
@@ 4.1 Evaluation Strategy
Self-awareness encompasses the abilities to recognize ‘knowns’ and ‘unknowns’. Accordingly, we introduce three metrics to measure a model’s self-awareness in the MM-SAP benchmark.
@@
Answer Rate Answer Acc Answer Rate Answer Acc Answer Rate
@@ 4.2 Inference Settings
For all the MLLMs tested in this study, we set the decoding temperature to $t=0$ and the decoding beam size to $b=1$. To reduce the uncertainty of the scores, each model is requested to predict the answer five times, with the order of the options randomly shuffled. We then calculate the mean of all scores as the result. More evaluation details are provided in Appendix B.
@@ 4.3 Main Results
A total of thirteen popular MLLMs were evaluated on our MM-SAP benchmark, including LLaVA-7B, LLaVA-13B Liu et al. (2023b, c), ShareGPT4V-7B, ShareGPT4V-13B Chen et al. (2023), CogVLM-17B Wang et al. (2023b), Yi-VL-6B Yi (2023), Qwen-VL-Chat, Qwen-VL-Plus, Qwen-VL-Max Bai et al. (2023), InfMLLM-7B Zhou et al. (2023), InternLM-XComposer2-VL-7B Dong et al. (2024), Gemini 1.0 Pro Vision Team et al. (2023), and GPT-4V OpenAI (2023b). The self-awareness scores $score_{sa}$ of these MLLMs are presented in Table 1.
@@ 4.4 Refusal Behavior of MLLMs
${\rm Unknown\ Knowns\ Rate}=$ (8)
@@ 6 Limitations
Acknowledgments
This work is supported by National Key R&D Program of China (No. 2022ZD0162101), STCSM (No. 21511101100, No. 22DZ2229005), and State Key Laboratory of UHD Video and Audio Production and Presentation.
Appendix A MM-SAP benchmark
We provide more details on the MM-SAP benchmark.
A.1 Statistic of Dataset
The average number of words in queries and options is presented in Table 5. Additionally, Table 6 shows the distribution of correct and refusal options.
average number of words
queries 9.8
options 2.8
Table 5: The average number of words in querie and options in MM-SAP.
A B C D E
correct options 20.3% 19.1% 18.8% 20.6% 21.2%
refusal options 21.6% 20.3% 18.7% 19.9% 19.5%
Table 6: The choice distribution of correct options and refusal options in MM-SAP.
A.2 The Categories of Questions in BeyondVisQA
@@
Appendix B Evaluation Detail
The prompts used for model evaluation are shown in Figure 6. To evaluate the responses, we employed two methods: calculating the perplexity (PPL) and directly matching the characters of the options. We applied these methods to LLaVA and ShareGPT4V as shown in table 7 and found that the results were nearly identical. Given that closed-source models like GPT-4V and Qwen-VL-Max cannot be evaluated using PPL calculations, we ultimately decided to evaluate answer correctness by directly matching the characters of the options for all models.
Figure 6: Prompts for model evaluation
Model BasicVisQA KnowVisQA BeyondVisQA Total
$score_{kk}$ $score_{kk}$ $score_{ku}$ $score_{ku}$ $score_{kk}$ $score_{ku}$ $score_{sa}$
LLaVA-7b(direct matching) 60.75 46.06 1.37 25.70 35.15 9.36 44.50
LLaVA-7b(PPL) 60.75 46.36 1.37 25.80 35.23 9.40 44.63
ShareGPT4V-7b(direct matching) 65.80 48.51 1.83 36.80 37.65 13.36 51.01
ShareGPT4V-7b(PPL) 65.75 48.71 1.83 37.05 37.68 13.45 51.14
Table 7: The results of LLaVA-7b and ShareGPT4V-7b with different evaluating method.
Appendix C Additional Examples in MM-SAP
In this section, we provide supplementary examples from our MM-SAP benchmark as shown in Figure 7, Figure 8, and Figure 9.